Rank the following times from smallest to largest:

* 4 nm/s

* 1 km/hr

*10 cm/s

*90 mm/s

*90 um/s
Answer #1

4 nm =4*10^-9 m/s

10cm= 1*10^-2m/s

90mm/s=1*10^-3m/s

90um/s =1*10^-6m/s

1km/hr=0.278m/s

So rank is :

4nm < 90um < 90mm < 10cm < 1km

(Rate please thanks )